OpenStudy (anonymous):

Frank has a batting average of 0.105. Write this decimal as a fraction in simplest form.

OpenStudy (cwrw238):

= 105 / 1000 can u simplify that?

OpenStudy (cwrw238):

hint; both numbers are divisible by 5

OpenStudy (anonymous):

im not getting it..

OpenStudy (cwrw238):

what don't you get ? 105/1000 ?

OpenStudy (anonymous):

its not one of my answers

OpenStudy (cwrw238):

right its not the fraction in its simplest form decimals are based on tenths, hundredths etc 0.2 = 2 /10 0.21 = 21/100 the 3rd number after decimal point is 1000th thats why 0.105 = 105/1000

OpenStudy (cwrw238):

now both 105 and 1000 are divisible by 5 105 / 5 = 21 1000 / 5 = 200 so in simplest form the fraction is 21/200 there are no common factors in 21 and 200

OpenStudy (cwrw238):

- you try to help some people and they just leave....